# Complementarity between Cosmic String Gravitational Waves and long lived particle searches in laboratory

### v1

@@ Abstract
Cosmic strings are powerful witnesses to cosmic events including any period of early matter domination. If such a period of matter domination was catalysed by metastable, long-lived particles, then there will be complementary signals to ascertain the nature of dark sector in experiments detecting primordial features in the gravitational wave (GW) power spectrum and laboratory searches for long-lived particles. We give explicit examples of global and local U(1) gauge extended dark sectors to demonstrate such a complementarity as the union of the two experiments reveals more information about the dark sector than either experiment. Demanding that Higgs-portal long-lived scalar be looked for, in various experiments such as DUNE, FASER, FASER-II, MATHUSLA, SHiP, we identify the parameter space which leads to complementary observables for GW detectors such as LISA and ET.
@@ I Introduction
Networks of cosmic strings (CS) are ubiquitous in cosmological models that include grand unification, as many symmetry breaking paths contain steps where the vacuum manifold has non-contractable loops Dror:2019syi ; Dunsky:2021tih ; King:2021gmj ; Fu:2023mdu . Such a network of CS continuously emits gravitational waves (GW) resulting in a spectrum that is flat during radiation domination, apart from the evolution of $g_{\ast}$ Cui:2018rwi . Any departure from such a flat spectrum speaks to a surprise in our cosmic history Cui:2018rwi ; Ghoshal:2023sfa ; Blasi:2020wpy ; Ferrer:2023uwz with a period of early matter domination providing a particularly striking feature in the otherwise flat spectrum Cui:2017ufi ; Ghoshal:2023sfa ; Datta:2020bht ; Samanta:2021zzk ; Chianese:2024gee ; Datta:2024bqp . Such a feature reveals two observables which yield information on the culprit for the early matter domination - the time matter domination began which is set by the highest frequency that departs from a flat spectrum and the duration of matter domination which corresponds to the width of the feature in frequency space. A period of early matter domination can be caused by a metastable, long-lived particle Coughlan:1983ci ; Starobinsky:1994bd ; Dine:1995uk ; Moroi:1999zb ; Ghoshal:2022ruy . Such a particle can arise in dark sectors as a mediator Ghoshal:2022ruy . The period of early matter domination arises when the metastable particle thermally decouples with an abundance much larger than the observed dark matter abundance. Since matter dilutes slower than radiation, as long as the particle is sufficiently long-lived there will be a period of matter domination.
Such long-lived particles are notoriously difficult to detect in traditional colliders due to their displaced vertex making them invisible Feng:2022inv . If the particle is sufficiently light, $\lesssim 0.3$ GeV there exist severe constraints from supernova cooling. For heavier particles, beam dump experiments and long-lived searches such as FASER, FASER-II FASER:2018bac ; FASER:2019aik , DUNE DUNE:2015lol and MATHUSLA Curtin:2018mvb ; Curtin:2023skh promise to probe large parts of the parameter space. However, any signal would leave the hidden sector woefully under-determined with only the lifetime and mass of the particle known. It is therefore crucial to find complementary sources of knowledge about such a long-lived particle. From primordial features in the stochastic gravitational wave background, we will show that one may obtain an independent measurement of the lifetime as well as a measurement of the annihilation cross-section between hidden and visible sectors. For the case of a hidden scalar singlet with three free parameters, one can test the BSM model due to the synergy between gravitational wave detectors and laboratory searches, for instance as proposed in Ref.Ghoshal:2022ruy in the context of inflationary GW. For generic studies on the impact of long-lived particles in inflationary tensor perturbations, see Ref.Berbig:2023yyy ; Borboruah:2024eha ; Borboruah:2024eal ; Ghoshal:2024gai ; Bernal:2020ywq ; Datta:2022tab ; Datta:2023vbs ; Chianese:2024nyw .
In this work, we consider the thermal history of a long-lived particle where CS are in the background bearing witness to periods of early matter domination in order to demonstrate the possibility of complementarity. In section II we give an explicit model. We then discuss the thermal history of our model in section III. Next, we give the imprint of CS onto the parameter space in section IV before discussing the ground-based signals in section V. Finally, we explicitly show the complementarity in section VI before concluding.
@@ II The model
Our main objective is to present a concrete scenario where complementary information about a model can be gleaned from ground-based experiments and the spectral shape of the GW background from CS, due to the presence of a particle sufficiently long-lived to catalyze a period of early matter domination. In order to realise such a non-standard cosmological evolution we consider the case where the Standard Model is extended by additional scalar fields. We also require a cosmic GW background generated by CS. Since a visible background requires a very high symmetry-breaking scale, the physics that produces such strings we assume to be weakly coupled from the additional scalars we will be discussing in detail. Keeping these objectives in view, the SM is extended by a $U(1)_{B-L}$ symmetry (with $B-L$ complex scalar $\Phi$) along with additional $\mathcal{Z}_{2}$-odd real scalar singlet, $S$. One can write the symmetry invariant full scalar potential as
@@
After breaking electroweak symmetry, $U(1)_{B-L}$ symmetry, and our global discrete $\mathcal{Z}_{2}$ symmetry, the scalar potential becomes a function of three classical fields $H=\frac{1}{\sqrt{2}}\begin{bmatrix}0\\ v+h\end{bmatrix}$, $S=(v_{s}+s)$, and $\Phi=\frac{1}{\sqrt{2}}(v_{BL}+\phi)$, $v$ being the EW scale, such that
@@
This of course is one corner of parameter space that features in our scan where we make no such restriction. Interestingly, the feeble effective coupling between the $B-L$ scalar and such a real scalar $S$ of the form
@@
can generate such a $B-L$ higgs dependent vev, $v_{s}=\sqrt{\lambda_{\Phi S}/\lambda_{S}}v_{BL}$ and following the electroweak symmetry breaking one can treat the model as a 2-field system having a mixed portal, controlled by the coupling $\lambda_{HS}$ thus providing an effective three-body interaction, as well as a mixing term
@@ III Cosmic history
The intermediate energy density of $S$ can be set by either freeze-out or freeze-in mechanisms, depending on the interaction rate $\langle\sigma_{S}v\rangle$. This decoupling leaves behind a residual energy density that redshifts as matter once $S$ becomes non-relativistic. This energy density can surpass that of radiation, leading to $S$ dominating if $\rho_{S}^{\rm FO/FI}(T)>\rho_{R}(T)$. For both decoupling mechanisms, we will need to calculate $H_{\rm EMD}$ as the time at which matter domination starts and ends will be the observables imprinted on the GW background by CS. For the freeze-out (FO) scenario, the redshifted frozen number density of $S$ particles at the beginning of the EMD is:
@@ IV Cosmic strings as a cosmic witness
Here the factor $0.1$ indicates that $90$ per cent of the loop population consists of small, highly-boosted loops. Consequently, these are usually red-shifted away without contributing to the GW radiation in any significant mannerBlanco-Pillado:2013qja . The loop formation efficiency factor $C_{\rm eff}$, depicting the decay rate at which local strings attain the asymptotic values or scaling solutions is given by $C_{\rm eff}\simeq 0.39$ and $5.4$ during the matter-dominated (MD) and radiation-dominated (RD) epochs respectively Gouttenoire:2019kij . For global strings, besides the loop formation, the long strings also lose energy into light Goldstone boson particle production, due to which the loop production efficiency becomes logarithmically suppressed. Analytically, this can be understood as $C_{\rm eff}\sim\mathcal{O}(1)$ across all cosmological epochs Chang:2019mza ; Gouttenoire:2019kij ; Chang:2021afa ; Ramberg:2019dgi . In order to accurately determine $C_{\rm eff}(t)$ one needs to solve the velocity-dependent one-scale (VOS) equations which govern the evolution of the string network. However, in this study, we employ simplistic scaling solutions instead, as they effectively capture the essential qualitative features of the model that we need to demonstrate for the complementarity studies that we allude to, in this paper.
@@
and since their formation at $\tilde{t}>t_{i}$, loops of length $l(\tilde{t})$ oscillate and radiate a discrete spectrum of GWs with frequencies set by the relation
@@
and $\Gamma=50$ for local cosmic strings Blanco-Pillado:2017oxo and global global strings Gorghetto:2021fsn . The value of $\delta$ depicts the scenario if this depends on whether high Fourier modes are dominated by cusps (for which $\delta=4/3$), kinks (for which $\delta=5/3$), or kink-kink collisions (for which $\delta=2$) Olmez:2010bi . This value greatly influences the frequency slope at the transition between early matter domination and radiation domination. Assuming $\delta=4/3$, that is the small-scale case to be dominated by cusps, we will proceed with the estimations. Since the loops continuously lose energy via the emission of GWs or via Goldstone boson production, their length $l$ keeps on shrinking as
@@
[…]
@@
where $\Omega_{r}h^{2}\simeq 4.2\times 10^{-5}$articleDataGroup:2020ssz and deviations from flatness may occur due to variations in the number of relativistic degrees of freedom accounted for in Gouttenoire:2019kij
@@ V Lab searches for long-lived particles
In this section we investigate the prospects of searching for this light scalar in intensity frontier and lifetime frontier experiments which look for light, weakly interacting, electrically neutral long-lived particles. These experiments are capable of probing up to extremely small mixing angles or couplings. Dark sectors with light degrees of freedom can be searched for, with a variety of experiments at the luminosity frontier, including proton Batell:2009di ; deNiverville:2011it ; deNiverville:2012ij ; Kahn:2014sra ; LBNE:2013dhi ; Soper:2014ska ; Dobrescu:2014ita ; Coloma:2015pih ; deNiverville:2016rqh ; MiniBooNE:2017nqe ; MiniBooNEDM:2018cxm ; MATHUSLA:2018bqv ; FASER:2018bac , electron Bjorken:2009mm ; Izaguirre:2013uxa ; Diamond:2013oda ; Izaguirre:2014dua ; Batell:2014mga ; BaBar:2017tiz ; Berlin:2018bsc ; Banerjee:2019pds and positron fixed target facilities that exist and upcoming all over the world Accardi:2020swt . Here we show, the allowed target parameter space of such dark sector lies well within the reach of several such experiments given the mass and mixing of the light scalar [^2] Light dark sector scenarios have been explored in great detail over the past few years in recent times Bramante:2016yju ; Alexander:2016aln ; Battaglieri:2017aum ; Darme:2017glc ; Winkler:2018qyg ; Egana-Ugrinovic:2019wzj ; Okada:2019opp ; Foroughi-Abari:2020gju ; Agrawal:2021dbo ; Nardi:2018cxi .. Such a tiny mixing angle is indeed within the reach of upcoming and current experiments like DUNE DUNE:2015lol ; Berryman:2019dme , FASER-II Feng:2017vli ; FASER:2018eoc ; FASER:2018bac ; FASER:2019aik , PS191 Bernardi:1985ny ; Gorbunov:2021ccu , DarkQuest-Phase2 Batell:2020vqn , MATHUSLA Curtin:2018mvb and SHiP SHiP:2015vad .
@@
$\Gamma_{H\rightarrow\ell\bar{\ell}}=\frac{m_{S}}{8\pi}\left(\frac{m_{f}}{v}\right)^{2}\left(1-\frac{4m_{f}^{2}}{m_{S}^{2}}\right)^{3/2},$ (31)
$\Gamma_{H\rightarrow q\bar{q}}=3\Gamma_{H\rightarrow\ell\bar{\ell}}\left[1+\frac{4\alpha}{3\pi}\left(\frac{9}{4}+\frac{3}{2}\log\frac{m_{f}^{2}}{m_{S}^{2}}\right)\right],$
@@
$\Gamma_{H\rightarrow KK}=\frac{m_{S}^{3}}{16\pi v^{2}}\left(\frac{2}{9}+\frac{11m_{\pi}^{2}}{m_{S}^{2}}\right)^{2}\left(1-\frac{4m_{\pi}^{2}}{m_{S}^{2}}\right)^{1/2},$
$\Gamma_{H\rightarrow\pi\pi}=\frac{m_{S}^{3}}{8\pi v^{2}}\frac{27}{13}\left(1-\frac{4m_{K}^{2}}{m_{S}^{2}}\right)^{3/2},$
Where $\Gamma_{\rm SM}^{\rm total}(m_{S})$ denotes the total decay width of the SM Higgs boson, restricted to channels that are kinematically open for the scalar decay and rescaled by the factor $m_{S}/m_{H}$. In the following sections, we will see that, since the decay width is proportional to $\sin^{2}\theta$, which can be very small, and as a result, the scalar can be quite long-lived to have the potential to imprint its characteristic signatures in lab searches.

### v2

# Complementarity between Cosmic String Gravitational Waves and long-lived particle searches in a laboratory

@@ Abstract
Cosmic strings are powerful witnesses to cosmic events including any period of early matter domination. If such a period of matter domination was catalysed by metastable, long-lived particles, then there will be complementary signals to ascertain the nature of dark sector in experiments detecting primordial features in the gravitational wave (GW) power spectrum and laboratory searches for long-lived particles. We give explicit examples of global and local U(1) gauge extended dark sectors to demonstrate such a complementarity as the union of the two experiments reveals more information about the dark sector than either experiment. Demanding that Higgs-portal long-lived scalar be looked for, in various experiments such as DUNE, FASER, FASER-II, MATHUSLA, and SHiP, we identify the parameter space that leads to complementary observables for GW detectors such as Laser Interferometer Space Antenna and Einstein Telescope.
@@ I Introduction
Networks of cosmic strings (CS) are ubiquitous in cosmological models that include grand unification, as many symmetry-breaking paths contain steps where the vacuum manifold has noncontractible loops Dror:2019syi; Dunsky:2021tih; King:2021gmj; Fu:2023mdu. Such a network of CS continuously emits gravitational waves (GW), resulting in a spectrum that is flat during radiation domination, apart from the evolution of $g_{\ast}$ Cui:2018rwi. Any departure from such a flat spectrum speaks to a surprise in our cosmic history Cui:2018rwi; Ghoshal:2023sfa; Blasi:2020wpy; Ferrer:2023uwz with a period of early matter domination providing a particularly striking feature in the otherwise flat spectrum Cui:2017ufi; Ghoshal:2023sfa; Datta:2020bht; Samanta:2021zzk; Chianese:2024gee; Datta:2024bqp. Such a feature reveals two observables which yield information on the culprit for the early matter domination - the time matter domination began, which is set by the highest frequency that departs from a flat spectrum, and the duration of matter domination, which corresponds to the width of the feature in frequency space. A period of early matter domination can be caused by a metastable, long-lived particle Coughlan:1983ci; Starobinsky:1994bd; Dine:1995uk; Moroi:1999zb; Ghoshal:2022ruy; Bishara:2024rtp. Such a particle can arise in dark sectors as a mediator Ghoshal:2022ruy. The period of early matter domination arises when the metastable particle thermally decouples with an abundance much larger than the observed dark matter abundance. Since matter dilutes slower than radiation, as long as the particle is sufficiently long-lived there will be a period of matter domination.
[…]
In this work, we consider the thermal history of a long-lived particle where CS are in the background bearing witness to periods of early matter domination in order to demonstrate the possibility of complementarity. In section II, we give an explicit model. We then discuss the thermal history of our model in section III. Next, we give the imprint of CS onto the parameter space in section IV before discussing the ground-based signals in section V. Finally, we explicitly show the complementarity in section VI before concluding.
@@ II The model
Our main objective is to present a concrete scenario where complementary information about a model can be gleaned from ground-based experiments and the spectral shape of the GW background from CS, due to the presence of a particle sufficiently long-lived to catalyze a period of early matter domination. In order to realise such a non-standard cosmological evolution we consider the case where the Standard Model (SM) is extended by additional scalar fields. We also require a cosmic GW background generated by CS. Since a visible background requires a very high symmetry-breaking scale, the physics that produces such strings we assume to be weakly coupled from the additional scalars we will be discussing in detail. Keeping these objectives in view, the SM is extended by a $U(1)_{B-L}$ symmetry (with BLB-L complex scalar $\Phi$) along with additional $\mathcal{Z}_{2}$-odd real scalar singlet, $S$. One can write the symmetry invariant full scalar potential as
@@
After breaking the electroweak (EW) symmetry, $U(1)_{B-L}$ symmetry, and our global discrete $\mathcal{Z}_{2}$ symmetry, the scalar potential becomes a function of three classical fields $H=\frac{1}{\sqrt{2}}\begin{bmatrix}0\\ v+h\end{bmatrix}$, $S=(v_{s}+s)$, and $\Phi=\frac{1}{\sqrt{2}}(v_{BL}+\phi)$, $v$ being the EW scale, such that
@@
The choice of this hierarchy among the VEV scales and masses imposes certain possible constraints on the potential parameters: specifically, we have $\lambda\gg\lambda_{s}\gg\lambda_{\Phi S}$, and the feeble mixing angle translates to $\lambda_{HS}\ll\lambda$. One motivation for such hierarchies $v_{BL}\gg v_{s}\gg v$ and $m_{S}\ll m_{H}$, can be thought to be related to neutrino mass generation mechanisms and dark matter physics, where $v_{BL}$ controls the seesaw scale Mohapatra:1980qe, which is typically very high for the vanilla type-I seesaw, where right-handed neutrinos have masses of the $\mathcal{O}(10^{12}$ GeV) or so. The scale $v_{s}$ is associated with dark sector, particularly in the context of freeze-in dark matter. Such a scalar whose search we propose in the complementary manner as long-lived particle then becomes a Higgs-portal like mediator for the dark matter, already discussed, see Refs. Ghoshal:2022ruy; Barman:2021lot. This of course is one corner of parameter space that features in our scan, where we make no such restriction. Interestingly, the feeble effective coupling between the BLB-L scalar and such a real scalar $S$ of the form
@@
can generate such a BLB-L Higgs dependent VEV, $v_{s}=\sqrt{\lambda_{\Phi S}/\lambda_{S}}v_{BL}$ and following the EW symmetry breaking one can treat the model as a 2-field system having a mixed portal, controlled by the coupling $\lambda_{HS}$ thus providing an effective three-body interaction, as well as a mixing term
@@ III Cosmic history
The intermediate energy density of $S$ can be set by either freeze-out or freeze-in mechanisms, depending on the interaction rate $\langle\sigma_{S}v\rangle$. This decoupling leaves behind a residual energy density that redshifts as matter once $S$ becomes non-relativistic. This energy density can surpass that of radiation, leading to $S$ dominating if $\rho_{S}^{\rm FO/FI}(T)>\rho_{R}(T)$. For both decoupling mechanisms, we need to calculate $H_{\rm EMD}$, as the times at which matter domination starts and ends will be the observables imprinted on the GW background by CS. For the freeze-out (FO) scenario, the redshifted frozen number density of $S$ particles at the beginning of the EMD is:
@@ IV Cosmic strings as a cosmic witness
Here the factor $0.1$ indicates that $90\%$ of the loop population consists of small, highly-boosted loops. Consequently, these are usually red-shifted away without contributing to the GW radiation in any significant mannerBlanco-Pillado:2013qja. The loop formation efficiency factor $C_{\rm eff}$, depicting the decay rate at which local strings attain the asymptotic values or scaling solutions is given by $C_{\rm eff}\simeq 0.39$ and $5.4$ during the matter-dominated (MD) and radiation-dominated (RD) epochs respectively Gouttenoire:2019kij. For global strings, besides the loop formation, the long strings also lose energy into light Goldstone boson particle production, due to which the loop production efficiency becomes logarithmically suppressed. Analytically, this can be understood as $C_{\rm eff}\sim\mathcal{O}(1)$ across all cosmological epochs Chang:2019mza; Gouttenoire:2019kij; Chang:2021afa; Ramberg:2019dgi. In order to accurately determine $C_{\rm eff}(t)$ one needs to solve the velocity-dependent one-scale (VOS) equations which govern the evolution of the string network. However, in this study, we employ simplistic scaling solutions instead, as they effectively capture the essential qualitative features of the model that we need to demonstrate for the complementarity studies that we allude to, in this paper.
@@
and since their formation at $\tilde{t}>t_{i}$, loops of length l(t~)l($\tilde${t}) oscillate and radiate a discrete spectrum of GWs with frequencies set by the relation
@@
and $\Gamma=50$ for local cosmic strings Blanco-Pillado:2017oxo and global strings Gorghetto:2021fsn. The value of $\delta$ depicts the scenario if this depends on whether high Fourier modes are dominated by cusps (for which $\delta=4/3$), kinks (for which $\delta=5/3$), or kink-kink collisions (for which $\delta=2$) Olmez:2010bi. This value greatly influences the frequency slope at the transition between early matter domination and radiation domination. Assuming $\delta=4/3$, that is the small-scale case to be dominated by cusps, we will proceed with the estimations. Since the loops continuously lose energy via the emission of GWs or via Goldstone boson production, their length $l$ keeps on shrinking as
@@
$\hskip 30.00005pt\int^{t_{0}}_{t_{\rm osc}}d\tilde{t}\,\frac{C_{\rm{eff}}(t_{i})}{t_{i}^{4}}\left[\frac{a(\tilde{t})}{a(t_{0})}\right]^{5}\left[\frac{a(t_{i})}{a(\tilde{t})}\right]^{3}\Theta\left(t_{i}-\frac{l_{*}}{\alpha}\right)\Theta(t_{i}-t_{\rm osc}).$ (22)
@@
where $\Omega_{r}h^{2}\simeq 4.2\times 10^{-5}$ParticleDataGroup:2020ssz and deviations from flatness may occur due to variations in the number of relativistic degrees of freedom accounted for in Gouttenoire:2019kij
@@ V Lab searches for long-lived particles
In this section we investigate the prospects of searching for this light scalar in intensity frontier and lifetime frontier experiments which look for light, weakly interacting, electrically neutral long-lived particles. These experiments are capable of probing up to extremely small mixing angles or couplings. Dark sectors with light degrees of freedom can be searched for, with a variety of experiments at the luminosity frontier, including proton Batell:2009di; deNiverville:2011it; deNiverville:2012ij; Kahn:2014sra; LBNE:2013dhi; Soper:2014ska; Dobrescu:2014ita; Coloma:2015pih; deNiverville:2016rqh; MiniBooNE:2017nqe; MiniBooNEDM:2018cxm; MATHUSLA:2018bqv; FASER:2018bac, electron Bjorken:2009mm; Izaguirre:2013uxa; Diamond:2013oda; Izaguirre:2014dua; Batell:2014mga; BaBar:2017tiz; Berlin:2018bsc; Banerjee:2019pds and positron fixed target facilities that exist and are being developed worldwide Accardi:2020swt. Here we show, the allowed target parameter space of such dark sector lies well within the reach of several such experiments given the mass and mixing of the light scalar [^3] Light dark sector scenarios have been explored in great detail over the past few years in recent times Bramante:2016yju; Alexander:2016aln; Battaglieri:2017aum; Darme:2017glc; Winkler:2018qyg; Egana-Ugrinovic:2019wzj; Okada:2019opp; Foroughi-Abari:2020gju; Agrawal:2021dbo; Nardi:2018cxi.. Such a tiny mixing angle is indeed within the reach of upcoming and current experiments like DUNE DUNE:2015lol; Berryman:2019dme, FASER-II Feng:2017vli; FASER:2018eoc; FASER:2018bac; FASER:2019aik, PS191 Bernardi:1985ny; Gorbunov:2021ccu, DarkQuest-Phase2 Batell:2020vqn, MATHUSLA Curtin:2018mvb and SHiP SHiP:2015vad.
@@
$\Gamma_{H\rightarrow\ell\bar{\ell}}=\frac{m_{S}}{8\pi}\left(\frac{m_{\ell}}{v}\right)^{2}\left(1-\frac{4m_{\ell}^{2}}{m_{S}^{2}}\right)^{3/2},$ (31)
$\Gamma_{H\rightarrow q\bar{q}}=3\Gamma_{H\rightarrow\ell\bar{\ell}}\bigg|_{\ell\rightarrow q}\left[1+\frac{4\alpha_{\rm QCD}}{3\pi}\left(\frac{9}{4}+\frac{3}{2}\log\frac{m_{q}^{2}}{m_{S}^{2}}\right)\right],$
@@
$\Gamma_{H\rightarrow\pi\pi}=\frac{m_{S}^{3}}{16\pi v^{2}}\left(\frac{2}{9}+\frac{11m_{\pi}^{2}}{m_{S}^{2}}\right)^{2}\left(1-\frac{4m_{\pi}^{2}}{m_{S}^{2}}\right)^{1/2},$
$\Gamma_{H\rightarrow KK}=\frac{m_{S}^{3}}{8\pi v^{2}}\frac{27}{13}\left(1-\frac{4m_{K}^{2}}{m_{S}^{2}}\right)^{3/2},$
where $\Gamma_{\rm SM}^{\rm total}(m_{S})$ denotes the total decay width of the SM Higgs boson, restricted to channels that are kinematically open for the scalar decay and rescaled by the factor $m_{S}/m_{H}$ and we have taken the QCD running coupling constant, denoted as $\alpha_{\rm QCD}$, to be approximately $\alpha_{\rm QCD}(M_{Z})\sim 0.1181$ [^4] Although, due to asymptotic freedom, $\alpha_{\rm QCD}$ decreases at high energies (short distances) and increases at low energies (long distances), we don’t expect much deviation in our results given the broad energy spectrum of interest for gravitational wave signatures.. In the following sections, we explore how the decay width, proportional to $\sin^{2}\theta$, can be very small, and as a result, the scalar can be quite long-lived to have the potential to imprint its characteristic signatures in lab searches.